OpenStudy (mica):

Prove: Sin3theta = 3sintheta - 4 sin^3theta I don't know where to start =/

OpenStudy (raden):

Sin3x = sin(x+2x) = sinxcos2x + cosxsin2x = sinx(1-2sin^2 x) + cosx(2sinxcosx) = sinx - 2sin^3 x + 2sinxcos^2 x = sinx - 2sin^3 x + 2sinx(1-sin^2 x) = sinx - 2sin^3 x + 2sinx -2sin^3 x = 3sinx -4sin^3 x